### v3

# Discovering Hierarchy-Grounded Domains with Adaptive Granularity for Clinical Domain Generalization

Abstract
Domain generalization has become a critical challenge in predictive healthcare, where different patient groups often exhibit shifting data distributions that degrade model performance. Still, regular domain generalization approaches often struggle in clinical settings due to (1) the absence of domain labels and (2) the lack of clinical insight integration. To address these challenges in healthcare, we aim to explore how medical ontologies can be used to discover dynamic yet hierarchy-grounded patient domains, a partitioning strategy that remains under-explored in prior work. Hence, we introduce UdonCare, a hierarchy-pruning method that iteratively divides patients into latent domains and retrieve domain-invariant (label) information from patient data. On two public datasets, UdonCare shows superiority over eight baselines across four representative clinical prediction tasks with substantial domain gaps, highlighting the potential of medical knowledge for enhancing model generalization.
Machine Learning, ICML
1 Introduction
The digitization of clinical data, notably electronic health records (EHR), has transformed healthcare by enabling efficient computational analysis. Current deep learning techniques have also achieved significant gains in diagnosis, mortality, and readmission prediction tasks (Poulain and Beheshti, 2024; Jiang et al., 2024). Still, these models trained on a set of training samples (source) often suffer performance drops when applied to a new test set (target) in a different domain. Domain shift denotes distributional changes between patient groups, such as data from different hospitals or time periods (Perone et al., 2019; Koh et al., 2021). Consequently, handling domain shift is a prerequisite for alleviating performance degradation in clinical predictive models (Yang et al., 2023a; Wu et al., 2023). It also aligns with the objective of most domain generalization (DG) methods, such as meta-learning (Balaji et al., 2018; Dou et al., 2019), feature alignment (Ganin et al., 2016; Li et al., 2018b), and latent-domain techniques (Matsuura and Harada, 2020; Wu et al., 2023).
In this work, we study DG problems in healthcare applications, which poses unique challenges beyond conventional DG scenarios: (1) Domain IDs (labels) are usually unseen in EHR datasets, but most DG solutions require the presence of domain IDs (Wu et al., 2023). Thus, some studies treat each patient as a unique domain (Dou et al., 2019; Yang et al., 2023a), which is overly fine-grained and unstable for model training, and others rely on broader partitions (e.g. institute & admission year), which overlook clinical heterogeneity (Zhang et al., 2021a; Guo et al., 2022). (2) Even though some DG methods do not rely on domain IDs (Arjovsky et al., 2019; Liu et al., 2021b), they overlook clinical semantics. For instance, prior DG methods (Matsuura and Harada, 2020; Wu et al., 2023) cluster instance features to form latent domains, but the resulting partitions are highly sensitive to training data. In practice, patient groups vary significantly, especially over longer admission periods. Existing DG methods can capture feature-level similarity, but not the progression of medical concepts. Hence, it is crucial to construct robust domains with explicit definitions grounded in clinical relevance.
To address these challenges, we explore the following research question: Instead of assuming the presence of domain IDs, can we leverage medical knowledge to guide models in discovering domains that are both adaptive and clinically meaningful? In hospitals, visiting patients are treated based on their medical history, which includes medical concepts such as diseases or medications. For instance, when dealing with heart failure patients, hospitals may categorize heart failure as a distinct domain or group it with other cardiovascular diseases. Similarly, heart failure corresponds to a leaf node under a higher-level node grouping cardiovascular diseases in ICD-9 ontologies (Organization and others, 1988), which motivates us to use a pruning algorithm to identify appropriate ancestor nodes for domain partitions.
However, translating medical ontologies into a form suitable for latent domain discovery is not a simple problem. Existing DG methods identify latent domains by clustering patient-level representations in vector spaces (Matsuura and Harada, 2020; Wu et al., 2023), which precludes the direct incorporation of ontology hierarchies. Meanwhile, medical ontologies are commonly used to enrich patient representations via graph neural networks (Jiang et al., 2024), but these ontologies are rarely explored as structural rules for discovering latent domains. Thus, we aim to bridge this disconnect and leverage ontologies not as auxiliary features, but as principled constraints that shape domain formation in a clinically meaningful way.
To this end, we introduce UdonCare, a framework that discovers Unseen domains with adaptive granularity to enhance model generalization in predictive healthcare. It ensures domain partitions remain consistent with clinical semantics while maintaining flexibility for adaptive generalization. Concretely, a pruning algorithm is developed to merges similar concepts on hierarchies and generate soft labels as domain IDs for patients. To explicitly remove covariates from patients, we propose a mutual learning network, which only retrieves label information upon the orthogonal factorization for final predictions. During training, model parameters are jointly optimized in an iterative manner across epochs, enabling the pruning module to adjust domain partitions based on the evolving latent embeddings. Our main contributions are enumerated as follows:
  • To the best of our knowledge, this is the first work using medical ontologies to tackle clinical DG problems. It reveals the potential of medical ontologies for discovering latent domains for handling covariates, rather than serving as feature enrichment.
  • UdonCare shows robustness across four predictive tasks on two public datasets, outperforming both clinical DG baselines (Yao et al., 2022; Wu et al., 2023) and regular DG baselines. For example, our model boosts the AUPRC score up to 11% over the best baselines.
  • We extend more analyses to show that UdonCare handles domain shifts through hierarchy-grounded domain partitions and accurate invariant feature learning, without sacrificing computational overhead.
2 Related Work
Domain Generalization (DG).  DG is pursuing adjusted models which are specially designed to remove domain-covariate features from hidden representation (Muandet et al., 2013). Most studies are dedicated to solve performance drops on target domain across diverse scenarios like image classification (Zhou et al., 2022; Ding et al., 2022), and they can be categorized as three different ways: (i) An intuitive way is to minimize the empirical source risk, either domain alignment (Ganin et al., 2016; Li et al., 2018b; Zhao et al., 2020) and invariant learning technique (Liu et al., 2021a; Zhang et al., 2022a; Wang et al., 2022) aim to convey little domain characteristics to acquire task-specific features. (ii) Contrastive learning (Kim et al., 2021; Jeon et al., 2021; Yao et al., 2022) becomes an alternative for data augmentation, studies employed the contrastive loss function to reduce the gap of representation distribution in one category. (iii) Meta-learning (Balaji et al., 2018; Dou et al., 2019) and ensemble-learning (Cha et al., 2021; Chu et al., 2022) approaches handle domain shifts through dynamic loss functions. However, they predefine either numerous or few domains in clinical settings (Wu et al., 2023), which motivates us to design a precise and efficient way to discover latent domains from learnable parameters.
DG in Clinical Prediction.  Prior works have shown that EHR predictive models often suffer performance drops when transferred to unseen patients with different data distributions (Perone et al., 2019; Koh et al., 2021). Some studies (Zhao et al., 2020; Zhang et al., 2021b) adopt feature adaptation to align domain shifts across multiple hospitals and different time periods (Reps et al., 2022; Zhang et al., 2022b). Recently, more works consider model generalization by mitigating the patient-specific domain shifts (Guo et al., 2022; Hai et al., 2024), which offers a flexible alternative in more scenarios. Moreover, ManyDG (Yang et al., 2023a) learns invariant features by treating each patient as unique domain; SLDG (Wu et al., 2023) develops a mixture-of-domain method to divide patient domains by concept-level features. However, they address domain partition with simplifying heuristics such as linear dependencies (Li et al., 2020), while support from medical knowledge beyond training data remains unexplored in clinical DG problems.
An additional discussion about ontology-based predictive models can be found in Appendix A.
Figure 1: The Overall Framework of UdonCare. The forward structure adds a domain pathway for mutual learning, extending beyond the backbone pathway of conventional predictive models. We feed patient data $\mathbf{x}$ into the backbone pathway, which learns patient features $\mathbf{p}$ through $f_{\phi,k}(\cdot)$ and produces the output prediction $\hat{y}_{p}$. In parallel, we obtain $\hat{y}_{h}$ along the domain pathway through $\mathrm{DiscoveryAlgo}(\cdot)$, $g_{\theta,k}(\cdot)$, and $h(\cdot)$. Here we can iteratively update domain partitions in $\mathbf{M}$ and update parameters on both pathways during training.
3 Preliminary
Clinical Prediction on EHR Data.   Each patient $\mathbf{x^{(i)}}$ consists of a longitudinal sequence of admissions $\{V^{(i)}_{1},V^{(i)}_{2},\dots,V^{(i)}_{T}\}$. We denote the entire set of medical concepts as $\mathcal{C}=\{c_{1},c_{2},...,c_{|\mathcal{C}|}\}$, and each admission contains a subset of $\mathcal{C}$. Most predictive models predict clinical outcomes $\mathbf{y}^{(i)}\in\{0,1\}^{d}$ at a future visit $V_{t+1}$, and they train a feature extractor $f_{\phi,k}(\cdot):\mathbf{x}_{k}\mapsto\mathbf{p}_{k}$ to encode $\mathbf{x}_{k}$ (omit $i$ in single patient) into patient embedding
$\mathbf{p}_{k}$ $=f_{\phi,k}(\mathbf{x}_{k})\in\mathbb{R}^{h},$ (1)
where $f_{\phi,k}$ is the encoder for feature key $k$, and $h$ denotes the dimension. When $f_{\phi}(\cdot)$ are trained on source training data sampled from distribution $P_{s}$, we often expect the learned model to also perform well on target data drawn from $P_{t}$ by assuming $P_{s}=P_{t}$. However, it is possible that these two distributions differs a lot $P_{s}\neq P_{t}$ due to spatial and temporal shifts, which motivates the need for DG solutions to factorize label information that remains invariant across clinical environments.
Concept-Specific Hierarchy.   In EHR data, certain medical concept $c_{i}\in\mathcal{C}$ always originates from a hierarchical encoding system, such as ICD (Organization and others, 1988) and ATC (Nahler and Nahler, 2009) codes. We define a concept-specific hierarchy $\mathcal{H}$ of $H$ levels, and denote $n^{(h)}_{i}$ as the $i$-th node on level $h$. Leaf nodes at level $H$ represent actual codes via the mapping $m:c_{i}\mapsto n^{(H)}_{i}$ with node feature $\mathbf{e}_{i}$ stored in $f_{\phi,k}$, and $\mathrm{Desc}(\cdot)$ denotes the set of descendant nodes. Note that, the root node $n^{(1)}_{1}$ at the top level subsumes all nodes in $\mathcal{H}$, and any two leaf nodes $n^{(H)}_{i}$ and $n^{(H)}_{j}$ share at least one common ancestor. In this paper, we focus on the disease hierarchy for method clarity, while more ontologies (treatments and medications) are explored in Appendix H.
4 Methodology
UdonCare iteratively operates two main steps (as shown in Figure 1): (1) developing a pruning algorithm for medical hierarchies to discover latent domains; (2) learning invariant features by factorizing patient embeddings in projection space. A notation table is also provided in Appendix C.
4.1 Hierarchy-Guided Domain Discovery
While domain IDs are unavailable in EHR data, it is intuitive that patients with similar medical diagnoses (concepts) often belong to the same domain. Naively partitioning patients by medical histories $\mathbf{x}$ implicitly induces soft domain labels, but leads to an exponential growth of latent domains with vocabulary size $|\mathcal{C}|$ (i.e., $2^{|\mathcal{C}|}$). To address this issue, we propose a hierarchy-guided domain discovery algorithm that prunes overly fine-grained nodes and assigns domain IDs at appropriate ancestor levels, forming a compact domain set that still covers all concepts (see Figure 2). An assignment matrix $\mathbf{M}$ can be generated to query domain IDs via
$\mathbf{M}:=\mathrm{DiscoveryAlgo}\Bigl(\{\mathbf{x}^{(i)}\}_{i=1}^{N_{\mathrm{tr}}}\Bigr)\in\{0,1\}^{N_{\mathrm{tr}}\times|\mathcal{C}^{{}^{\prime}}|},$ (2)
where $N_{\mathrm{tr}}$ is the number of training patients, $|\mathcal{C^{{}^{\prime}}}|$ is the pruned vocabulary size. It merges fine-grained codes into higher-level clusters, allowing patients with or without a particular disease to occupy different latent domains as needed.
Initialization on Domain IDs. Following the previous settings (Lu et al., 2021; Jiang et al., 2024), we only focus on patients with multiple admissions $T\geq 2$. This sequence should be then converted into a single vector that consolidates all prior visits. Hence, we aggregate patient data $\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}\}$ into a unified representation $\mathbf{X}=\bigvee_{t=1}^{T}\mathbf{x}_{t}$, where medical concepts shown in each admission are merged to form a accumulated medical history. Note that, $\mathbf{X}$ is the most fine-grained domain partition.
Initialization on Node Features. With $\mathbf{X}$ for each patient, concept-specific ontology $\mathcal{H}$ with node features are required to decide whether fine-grained medical concepts group diseases into higher-level clusters or are preserved. We initialize leaf-node features $\mathbf{e}_{i}\in\mathbb{R}^{h}$ by: (1) for present code $c_{i}$ in $\mathcal{S}$, $\mathbf{e}_{i}$ is initialized from embedding layer $\text{E}(e_{1},\dots,e_{|\mathcal{C}|})$ in $f_{\phi,k}(\cdot)$; (2) for absent code $c_{i}$ in $\mathcal{S}$, $\mathbf{e}_{i}$ is its embedding of entity name through ClinicalBERT (Huang et al., 2019). The feature of a parent node $\mathbf{e}^{(h-1)}_{n_{i}}$ is then computed from the embeddings of its descendants at level $h$:
$\mathbf{e}^{(h-1)}_{i}\;:=\;\frac{1}{|\mathrm{Desc}(n^{(h-1)}_{i})|}\sum_{n\in\mathrm{Desc}(n_{i})}\mathbf{e}^{(h)}_{n}$, (3)
which extends $\text{E}(e_{1},\dots,e_{|\mathcal{C}|})$ to $\text{E}(e_{1},\dots,e_{|\mathcal{H}|})$ over the entire hierarchy. Still, it fails to capture the hierarchical distances in $\mathcal{H}$. For example, two leaf nodes might have similar features without sharing the same parent node. Under this scenario, their lowest common ancestor (LCA) $n_{\text{LCA}}(d_{i},d_{j})$ could be an alternative reflecting such similarity. Hence, we mimic the principle of hierarchical clustering (Johnson, 1967) in this stage. Concretely, we determine the most similar node pair $(d_{i},d_{j})$ based on their cosine similarity $\text{cos}(\mathbf{e}_{d_{i}},\mathbf{e}_{d_{j}})$, and then update their LCA embedding
$\mathbf{e}_{\mathrm{LCA}(d_{i},d_{j})}\leftarrow\mathrm{Average}\bigl(\mathbf{e}_{\mathrm{LCA}(d_{i},d_{j})},\mathbf{e}_{d_{i}},\mathbf{e}_{d_{j}}\bigr).$ (4)
This process continues down the similarity list until the remaining highest similarity falls below the threshold $\rho=0.3$. See Appendix J for details of the information flow.
Figure 2: An illustration of hierarchy-guided domain discovery.
Node Scoring. For each node $n\in\mathcal{H}$, we define $S(n)$ to identify which node is a good “candidate” for final selection. Motivated by the information gains (Song and Ying, 2015), $S(n)$ involves three indicators (see in Appendix K), coverage $\mathrm{cov}(n)$, purity $\mathrm{pur}(n)$, and depth $\mathrm{dep}(n)$, via
$S(n)=\exp(\mathrm{pur}(n))+\alpha\cdot\bigl(\mathrm{cov}(n)\times\mathrm{dep}(n)\bigr)$
$=\exp\Bigl(\mathbb{E}_{m\in\mathcal{M}}[\mathrm{sim}(\mathbf{e}_{n},\mathbf{e}_{m})]\Bigr)+\alpha\cdot(\frac{|\mathcal{M}|}{|\mathcal{L}|}\cdot\frac{h}{H}),$ (5)
where $\mathcal{M}$ is equivalent to Desc(n); $\mathbb{E}(\cdot)$ denotes the mathematical expectation; $\alpha$ and $\exp(\cdot)$ act as scaling factors, which regularize the selection avoiding either too low or high level. Consequently, the score matrix $S(s_{1},\dots,s_{|\mathcal{H}|})$ is obtained after scoring all nodes in the hierarchy.
Hierarchy Pruning. Once $S(n)$ is computed, we perform a bottom-up pass over $\mathcal{H}$ to generate a candidate set of pruned nodes. Let $p$ be the parent node with children $\{c_{1},\dots,c_{r}\}$. There are three possible situations upon comparing score $S(p)$ with its children scores $\{S(c_{i})\}^{r}$:
Case 1. If $S(p)>\max(\{S(c_{i})\}^{r})$, we include parent node $p$ and exclude its children.
Case 2. If $S(p)<\min(\{S(c_{i})\}^{r})$, we discard parent node $p$ and select all children.
Case 3. Otherwise, we tentatively include $p$ but mark it for further resolution in the next step.
Note that, if $p$ is discarded in the above cases, $\{c_{1},\dots,c_{r}\}$ are promoted to become the children of $p$’s parent. After the first iteration on the score matrix, a candidate subset $\mathcal{C}_{0}$ is obtained; however, it remains sub-optimal and requires further evaluation. To this end, we construct a list of tuples $A$ of length $N$, where each element $A[n]:=(p,{c_{1},\dots,c_{r}})$ represents a flagged parent–child pair.
Domain Searching. Given $N$ flagged pairs, the complexity of searching approaches grows exponentially $O(2^{N})$. Inspired by the Beam-Search algorithm (Lowerre, 1976), we get decisions within the top-k pruning candidates $O(kN)$. For each flagged pair, we either (1) unify them by including the parent or (2) retain the children as distinct pruned nodes. Those pairs that have not yet been decided default to retaining parent node to ensure complete coverage. We adopt the Silhouette Score (Shahapure and Nicholas, 2020) as the evaluation metric. Specifically, UdonCare first apply UMAP (McInnes et al., 2018) on leaf node features for dimensionality reduction, and it then calculate the score on this full hierarchical partition, reflecting the global separation quality under the current pruning scheme. This process iterates over all flagged node pairs in order by updating pruning subsets in $\{\mathcal{C}_{1},\mathcal{C}_{2},\dots,\mathcal{C}_{N}\}$ to make selection with the optimal evaluation result.
Domain Decision. With the pruned selection $\mathcal{C}_{N}$ ($\mathcal{C}^{\prime}$ in Eq. (2)), we update $\mathbf{M}$ by converting patient data into these pruned clusters in $\mathcal{C}_{N}$. If patient $\mathbf{x}^{(i)}$ has at least one leaf code $d$ that descends from the pruned node, we update $\mathbf{M}[i,j]=1$; otherwise, $\mathbf{M}[i,j]=0$. Consequently, each row $\mathbf{M}[i,:]$ represents the domain labels of the $i$-th patient, thereby defining a structured domain partitions over patients.
4.2 Mutual Forward Learning
Each domain can be viewed as the latent representation $\mathbf{r}$ sampled from a meta domain distribution $p(\cdot)$, so that we can identify $\mathbf{r}$ and then factorize $p(y|\mathbf{x})$ into $\int p(y|\mathbf{x},\mathbf{r})p(\mathbf{r}|\mathbf{x})\,d\mathbf{r}$ by approximating $q(\mathbf{r})\sim p(\mathbf{r}|\mathbf{x})$ given data samples $\mathbf{x}$. Subsequently, a domain encoder $p(\mathbf{r}|\mathbf{x})$ and a label predictor $p(y|\mathbf{x},\mathbf{r})$ are needed for inference. Here we parameterize the domain encoder $p(\mathbf{r}|\mathbf{x})$ as a network $g_{\theta}(\cdot)$ with parameter $\theta$. Since the pruned output matrix $\mathbf{M}$ (see Section 4.1) maps each training sample $\mathbf{x}$ to $\mathbf{m}$ (soft-label domain IDs), we apply $g_{\theta}(\cdot)$ to $\mathbf{m}$ to estimate the domain factor $\mathbf{r}:=g_{\theta}(\mathbf{m})$. Although $\mathbf{r}$ represents a probabilistic domain variable, we implement $g_{\theta}$ as a deterministic Multi-Layer Perceptron (MLP) for the prediction task. Next, we compute invariant features using a non-parametric function $h(\cdot):(\mathbf{r},\mathbf{p})\mapsto\mathbf{h}$, which fuses $\mathbf{r}$ (domain-level representation) and $\mathbf{p}$ (patient-level representation) as input features for the label predictor $p(y|\mathbf{x},\mathbf{r})$.
Self-Supervised Domain Encoder. The main concerns on training domain encoder is how to ensure $g_{\theta}(\cdot)$ can extract valid domain information from patients, which is ignored by some works (Finn et al., 2017; Li et al., 2018a; Yang et al., 2023a). A regulation method is then developed during the encoder training phase. Concretely, pseudo domain labels $\mathbf{m}$ help us divide patients into latent domains, where averaging patient-specific features $\bar{\mathbf{p}}$ could provide guidance for $g_{\theta}(\cdot)$ in identifying domain information. Hence, we adopt a pretraining task and update $\theta$ based on patient embeddings $\mathbf{p}$ from $f_{\phi}(\cdot)$ by minimizing loss function
$\mathcal{L}_{r}[g_{\theta}(\mathbf{m}),\bar{\mathbf{p}}]:=\mathrm{MSE}(\mathbf{r},\;\mathbb{E}[\mathbf{p}|\mathbf{m}])+\frac{\|\mathbf{r}_{\mu}-\mathbf{p}_{\mu}\|_{\mathcal{F}}^{2}}{\|\mathbf{p}_{\mu}\|_{\mathcal{F}}^{2}},$ (6)
where $\mathbb{E}[\mathbf{p}|\mathbf{m}]$ denotes the average embedding associated with domain IDs, and $\|\mathbf{r}_{\mu}-\mathbf{p}_{\mu}\|_{\mathcal{F}}^{2}$ measures the Maximum Mean Discrepancy (Borgwardt et al., 2006) with the norm $\mathcal{F}$ to reduce distributional gaps. The subscript $\mu$ indicates batch-level averages. The domain encoder $g_{\theta}(\cdot)$ can then approximate domain features $\mathbf{r}$ through both patient-level inputs $\mathbf{p}$ and $\mathbf{m}$.
Invariant Feature Projection Learning. In Eq. (6), both $\mathbf{r}$ and $\mathbf{p}$ are rescaled into a shared vector space with comparable magnitudes. Hence, we can directly apply an orthogonal projection approach (as in early studies (Bousmalis et al., 2016; Shen et al., 2022; Yang et al., 2023a)) to obtain the invariant feature $\mathbf{h}$ by subtracting the parallel component of $\mathbf{p}$ in this shared vector space. We formalize this in $h(\cdot)$ as shown in Eq. (7):
$\mathbf{h}:=\mathbf{p}-\tilde{\mathbf{r}},\;\text{where}\;\tilde{\mathbf{r}}=\mathbf{r}\cdot\langle\frac{\mathbf{p}}{\|\mathbf{r}\|},\frac{\mathbf{r}}{\|\mathbf{r}\|}\rangle.$ (7)
Here, $\tilde{\mathbf{r}}$ is the component of $\mathbf{p}$ that is parallel to $\mathbf{r}$ with domain covariates, while $\mathbf{h}$ is the remainder and thus invariant to domain shifts. We thus obtain invariant features $\mathbf{h}$ without additional parameters, and $h(\cdot)$ serves as an essential pre-processing step before making prediction.
4.3 Training and Inference
Iterative Training. To train UdonCare, we feed each data sample $\mathbf{x}$ into the hierarchy-pruning module to obtain its latent domain $\mathbf{m}$, and then perform two cross-reference steps under a mutual learning architecture. Rather than updating the model continuously in each epoch, we adopt an iterative training strategy, which prior studies (Cui et al., 2019; Sofiiuk et al., 2022) have shown can reduce training time while maintaining comparable predictive performance. [^1] For example, we set iterations $I=3$ and epochs $N=100$ by obtaining pretrained parameters for 40 epochs and then updating $\mathbf{M}$ iteratively every 20 epochs, yielding a total of 100 epochs. We iteratively update the model weights and regenerate domain assignments every 20 epochs in our experiment. Before each iteration, we reinitialize the parameters in $g_{\theta}(\cdot)$, because the input shape of $\mathbf{m}$ may change due to updated code-level embeddings. We also provide the pseudo-code of UdonCare in Appendix B.
Mutual Inference. After the orthogonal projection, we apply the network $q_{\xi}(\cdot)$ (operates on space $\mathbf{p}$) as a post-step to parameterize the label predictor $p(y|\mathbf{x},\mathbf{r})$; We can also parameterize $p(y|\mathbf{x})$ through the regular decoder network $d_{\eta}(\cdot)$ fed by patient embeddings from backbone $f_{\phi}(\cdot)$.
$p(y|\mathbf{x})\sim\hat{y}_{p}$ $=d_{\eta}(\mathbf{p})=d_{\eta}(f_{\phi}(\mathbf{x})),$
$p(y|\mathbf{x},\mathbf{r})\sim\hat{y}_{h}$ $=q_{\xi}(\mathbf{h})=q_{\xi}(h(g_{\theta}(\mathbf{m}),\mathbf{p})).$ (8)
Both $q_{\xi}(\cdot)$ and $d_{\eta}(\cdot)$ are linear classifiers with learnable parameter matrices. A loss function $\mathcal{L}$ is then applied to add label supervision for downstream predictive tasks. We integrate these two predictors into a collaborative framework, with the mutual inference objective:
[…]
5 Experiments
5.1 Experiment Setups
Table 1: Performance comparison of four prediction tasks on MIMIC-III/MIMIC-IV. We report the average performance (%) and the standard deviation (in bracket) over 5 runs. We also bold for the best and underscore for the second best model.
Model Task 1: Mortality Prediction Task 2: Readmission Prediction
MIMIC-III MIMIC-IV MIMIC-III MIMIC-IV
AUPRC AUROC AUPRC AUROC AUPRC AUROC AUPRC AUROC
Oracle 16.95 (0.48) 70.61 (0.58) 8.46 (0.53) 68.92 (0.47) 74.42 (0.43) 69.74 (0.48) 69.37 (0.12) 68.47 (0.11)
Base 11.05 (0.65) 55.01 (0.98) 3.97 (0.47) 59.13 (0.76) 63.45 (0.84) 62.15 (0.73) 58.46 (0.24) 60.56 (0.31)
DANN 12.80 (0.52) 63.29 (0.69) 4.41 (0.45) 63.82 (0.46) 64.02 (0.61) 63.89 (0.75) 60.89 (0.11) 61.22 (0.23)
CondAdv 13.60 (0.52) 65.22 (0.53) 5.65 (0.62) 64.27 (0.68) 67.72 (0.41) 65.47 (0.50) 62.78 (0.14) 62.29 (0.19)
MLDG 13.25 (0.44) 64.07 (0.63) 4.75 (0.38) 62.75 (0.57) 66.00 (0.55) 65.59 (0.59) 60.20 (0.23) 62.24 (0.28)
IRM 13.50 (0.48) 65.41 (0.54) 4.14 (0.52) 62.36 (0.61) 67.07 (0.43) 65.93 (0.55) 61.81 (0.14) 61.84 (0.17)
PCL 13.70 (0.49) 64.55 (0.61) 5.35 (0.49) 64.70 (0.55) 65.03 (0.49) 66.36 (0.59) 61.03 (0.28) 63.04 (0.26)
ManyDG 14.10 (0.54) 66.27 (0.52) 6.06 (0.31) 64.66 (0.32) 68.63 (0.45) 67.03 (0.56) 63.48 (0.25) 64.07 (0.24)
SLDG 13.20 (0.47) 63.63 (0.56) 4.58 (0.44) 63.24 (0.60) 65.73 (0.53) 65.80 (0.50) 61.55 (0.14) 65.29 (0.16)
UdonCare 15.90 (0.36) 68.78 (0.40) 6.81 (0.27) 66.73 (0.48) 71.17 (0.32) 67.68 (0.36) 65.61 (0.10) 67.26 (0.25)
Model Task 3: Drug Recommendation Task 4: Diagnosis Prediction
MIMIC-III MIMIC-IV MIMIC-III MIMIC-IV
AUPRC F1-score AUPRC F1-score w-F1 R@10 w-F1 R@10
Oracle 80.43 (0.14) 67.41 (0.29) 74.31 (0.25) 61.28 (0.22) 26.97 (0.13) 39.46 (0.27) 28.12 (0.11) 40.53 (0.16)
Base 68.83 (0.19) 57.31 (0.35) 66.94 (0.18) 53.13 (0.18) 21.32 (0.19) 31.08 (0.23) 20.07 (0.12) 31.52 (0.18)
DANN 75.18 (0.15) 60.55 (0.30) 69.63 (0.27) 53.43 (0.26) 22.08 (0.15) 34.27 (0.28) 24.05 (0.14) 35.21 (0.20)
CondAdv 76.59 (0.18) 64.42 (0.25) 71.48 (0.15) 55.62 (0.29) 23.09 (0.18) 36.27 (0.21) 26.13 (0.12) 37.35 (0.18)
MLDG 74.76 (0.17) 58.85 (0.33) 70.29 (0.27) 56.77 (0.16) 21.34 (0.17) 34.16 (0.26) 24.17 (0.13) 34.72 (0.19)
IRM 69.47 (0.16) 62.20 (0.30) 69.12 (0.14) 54.57 (0.18) 22.27 (0.16) 32.86 (0.23) 23.54 (0.15) 34.12 (0.21)
ManyDG 77.29 (0.14) 63.70 (0.28) 71.26 (0.19) 55.27 (0.19) 23.39 (0.14) 36.49 (0.27) 25.91 (0.13) 37.04 (0.20)
UdonCare 78.16 (0.18) 66.70 (0.29) 73.07 (0.33) 59.23 (0.14) 24.11 (0.18) 38.37 (0.26) 27.31 (0.11) 39.41 (0.17)
Predictive Tasks. We evaluate our approach on four representative tasks: (1) Mortality Prediction, which determines whether a patient will pass away by a specified time horizon after discharge. This is a binary classification task. (2) Readmission Prediction, which checks if a patient will be readmitted within a predefined window (e.g., next 15 days) following discharge. This is also framed as a binary classification. (3) Diagnosis Prediction, which forecasts the set of diagnoses (ICD-9-CM codes) for the patient’s next hospital visit based on prior visits. This requires multi-label classification. (4) Drug Recommendation, which suggests a set of medications (ATC-4 codes (Nahler and Nahler, 2009)) for the upcoming visit, also formulated as multi-label classification. These tasks reflect diverse clinical needs and provide a rigorous benchmark for evaluating DG methods.
Evaluation Metrics. Both readmission and mortality prediction are binary classification tasks, we calculate the Area Under the Precision-Recall Curve (AUPRC) and the Area Under the Receiver Operating Characteristic Curve (AUROC) scores due to the imbalanced label distribution. For drug recommendation, we evaluate predictions of all DG approaches by AUPRC and F1 scores, the same setting as ManyDG (Yang et al., 2023a) (i.e. $d<120$). For diagnosis prediction, we decide accurate prediction by weighted $\text{F}_{1}$ score as in Timeline and top-10 recall as in DoctorAI (Choi et al., 2016), since the former one measures the overall prediction on all classes (i.e. $d>4500$) and the latter one have concentration on positive code with low frequency.
Table 2: Cosine similarity of linear weights on $\mathbf{p}$, $\widetilde{\mathbf{r}}$, and $\mathbf{h}$. Cosine similarity is averaged across classes and evaluated over 3 runs.
Cosine Similarity Mortality Prediction Readmission Prediction Drug Recommendation Diagnosis Prediction
$W_{d}(\mathbf{p}\rightarrow\text{labels})$ vs. $W_{d}(\mathbf{h}\rightarrow\text{labels})$ $0.7831\pm 0.0174$ $0.4813\pm 0.0391$ $0.6546\pm 0.0237$ $0.8654\pm 0.0198$
$W_{d}(\mathbf{p}\rightarrow\text{domains})$ vs. $W_{d}(\widetilde{\mathbf{r}}\rightarrow\text{domains})$ $0.3427\pm 0.0088$ $0.1957\pm 0.0314$ $0.2753\pm 0.0274$ $0.2133\pm 0.0036$
$W_{d}(\mathbf{p}\rightarrow\text{labels})$ vs. $W_{d}(\mathbf{p}\rightarrow\text{domains})$ $1239\pm 0.0126 0.0$$0794\pm 0.0392$ $1251\pm 0.0212 0.0$$0972\pm 0.0095$
* $W_{\text{d}}(\cdot)$ represents the learned linear weights. For example, $W_{\text{d}}(\mathbf{p}\rightarrow\text{domains})$ denotes training a linear model on $\mathbf{p}$ to predict domains.
Datasets & Data Split. We conduct experiments on two publicly available EHR databases, MIMIC-III and MIMIC-IV, which are widely used in clinical prediction (Johnson et al., 2016, 2023). MIMIC-III covers ICU admissions from 2001 to 2012, while MIMIC-IV spans 2008 to 2019. To avoid overlapping time ranges with MIMIC-III, we only retain patients from the years 2013201320192019 in MIMIC-IV. For each set of experiments, we extract 6,4976{,}497 and 49,55849{,}558 patients with multiple visits ($T\geq 2$) from both datasets as shown in Appendix E Table 5. Different from random data splitting, we evaluate our model under domain shifts induced by demographic and temporal factors. For the MIMIC-III dataset, patients with Medicare or Medicaid insurance are used as the source domain, while those with private insurance are treated as the target domain. Following the setting in SLDG (Wu et al., 2023) on the MIMIC-IV dataset, patients admitted after 2017 are assigned to the target domain, while all preceding patients are used as the source domain. We further divide the target subset into validation and test subsets using a fixed ratio of 20%:80%. Note that, to prevent information leakage, there is no patient overlap between the training and test sets.
Baselines. We compare UdonCare with two naïve baselines, five general DG baselines and two most recent clinical DG baselines: (1) Naïve Baselines: Oracle, trained backbone encoder directly on the target data, and Base, trained solely on the source data. The difference between metrics from Oracle (upper bound) and Base (lower bound) can show the distribution gaps between source and target data. (2) Typical DG Baselines: DANN (Ganin et al., 2016), CondAdv (Isola et al., 2017), MLDG (Li et al., 2018a), IRM (Arjovsky et al., 2019), and PCL (Yao et al., 2022). (3) Clinical DG Baselines: ManyDG (Yang et al., 2023a), and SLDG (Wu et al., 2023). Since drug recommendation and diagnosis prediction are multi-label classification, we drop PCL and SLDG in these two tasks due to their setting limitation. Note that, all baselines follow the same source/target definition as in the Data Split section, and the test set remains unseen to all models except Oracle. The results reported in following sections are the performance evaluated on the test set. More details of the baselines can be found in Appendix D, and UdonCare is implemented as described in Appendix F.
5.2 Main Results
Table 1 presents results on four classification tasks using MIMIC-III and MIMIC-IV. First, the performance gap between the Oracle and Base methods is substantial, showing the presence of considerable domain differences. DANN (Ganin et al., 2016) and MLDG (Li et al., 2018a), both relying on coarse domain partitions, show minimal improvements, likely due to difficulties in extracting consistent features from coarse partitions. PCL (Yao et al., 2022) exhibits a slight gain through proxy-to-sample relationships. Meanwhile, IRM (Arjovsky et al., 2019) and CondAdv (Isola et al., 2017) perform better by incorporating regularization or recurrent structures for predictions. Among clinical-specific baselines, ManyDG (Yang et al., 2023a) achieves the best results by leveraging mutual reconstruction, and SLDG (Wu et al., 2023) sees only modest improvement due to its reliance on the most recent admissions. Lastly, UdonCare surpasses over all baselines across all tasks. Specifically, UdonCare achieves a relative AUPRC improvement of around 11% (12%) for mortality prediction in MIMIC-III (MIMIC-IV), and 4% (3%) for readmission prediction on the MIMIC-III (MIMIC-IV) dataset.
Beyond the above results, we further extend our experiments by using GAMENet (Shang et al., 2019b) and CGL (Lu et al., 2021) as our backbone, and evaluate model generalization across hospitals on the eICU dataset (see Appendix G).
Figure 3: Domains among diseases of the circulatory system identified by UdonCare. The left table shows how the discovered domains correspond to specific ICD-9-CM diseases (codes) in the hierarchy, while the right figure is the scatter plot using t-SNE upon code-level embeddings.
5.3 Analysis of Discovered Domain
To illustrate how UdonCare behaves, we evaluate discovered domains on diagnosis prediction using MIMIC-III. Here, we focus on three kinds of heart related conditions from diseases of the circulatory system: heart failure (HF, 15 types), essential hypertension (EH. 3 types), and acute rheumatic fever (ARF, 7 types). In the left table of Figure 3, we can observe that the mid-level “heart failure”, high-level “hypertensive disease”, and several leaf-level nodes are chosen to represent these subtypes. These decisions reflect the same patterns of code-embedding geometry (right scatterplot plot): HF codes form a compact cluster, EH codes cluster even more tightly and lie near the HF group, while ARF codes are widely dispersed. Moreover, it is also consistent with medical knowledge: HF and EH are clinically cohesive, while ARF shows greater heterogeneity due to variable cardiac involvement. Thus, these results align with both data-driven and clinical logic.
5.4 More Quantitative Analyses
Effectiveness of Decomposition. Following the setting of (Shen et al., 2022; Yang et al., 2023a), a linear classifier $d$ can be trained on the embedding to predict either (i) labels or (ii) domains. After training such a predictor, the cosine similarity can be calculated in terms of the learned weights to quantify the feature dimension overlaps. Note that $\mathbf{p}$, $\widetilde{\mathbf{r}}$, and $\mathbf{h}$ are normalized dimension-wise to ensure that each dimension is comparable. The results are shown in Table 2. In general, the third row has lower cosine similarities than the first two rows, which indicates that there is mostly non-overlap between feature dimensions predicting labels and domains. Moreover, the first two rows give relatively higher similarity and imply the domain and label information are separated from $\mathbf{p}$ into domain features $\tilde{\mathbf{r}}$ (scaling from $\mathbf{r}$) and invariant features $\mathbf{h}$. It provides the quantitative evidence that UdonCare stores domain and label information along distinct dimensions. We provide more explanations in Appendix M. Under the same setting as Table 2, we also attach the convergence process in detail (see Appendix I).
Figure 4: Effectiveness of Domain Discovery. The left figure shows the effect of the number of iteration on AUPRC, and the right one shows comparison among variants upon UdonCare.
Ablation Analysis. We evaluate whether the designed domain-discovery algorithm is effective for prediction. Given a lookup embedding table for condition concepts, we need to group similar codes to reduce dimensionality. Hence, (a) k-Means clustering, (b) hierarchical clustering, and (c) tree pruning of the information gain algorithms can be adopted to simplify the process. We use drug recommendation tasks as an example to assess performance. The results can be found in Figure 4. The left figure illustrates the trade-off between accuracy and computing time at three iterations. We observe that k-Means performs the worst, largely because of its limitations in determining the optimal number of clusters via grid search. Hierarchical clustering performs better than k-Means but lacks structured guidance from the medical hierarchy. Tree pruning outperforms both methods by leveraging medical ontologies, demonstrating the importance of knowledge-driven clustering. Moreover, UdonCare outperforms all these methods by incorporating more precise domain IDs through iterative beam-search updates. These results highlight the critical role of medical ontologies in domain discovery and the advantages of adaptive refinement for learning meaningful structures.
Table 3: Running Time Comparison of Drug Recommendation (seconds per epoch). Note that SLDG uses only the most recent admissions for prediction, following the original setting.
Model MIMIC-III MIMIC-IV
Base 3.206 $\pm$ 0.1219 6.943 $\pm$ 0.2342
ManyDG 5.462 $\pm$ 0.2648 9.215 $\pm$ 0.3781
SLDG 4.518 $\pm$ 0.0256 8.439 $\pm$ 0.1329
UdonCare 4.320 $\pm$ 0.1473 7.871 $\pm$ 0.2415
Runtime Analysis Next, we compare the training time of UdonCare with two other clinical DG baselines, ManyDG and SLDG, as shown in Table 3. All runtimes are measured on an NVIDIA L40S GPU. We find that an iterative training strategy effectively balances computational overhead and performance for the entire framework. We observe that UdonCare requires a training time comparable to SLDG, since both rely on iterative parameter updates that reduce the frequency of model adjustments during inference. ManyDG consumes more time than others, primarily because its domain assumption spawns numerous latent domains for subsequent computations.
Figure 5: Performance by Training Size. We evaluate drug recommendation on MIMIC-IV, and the x-axis indicate % of the entire training data. The dotted lines divide [1, 10] and [10, 100].
Effect of Training Data Size. Lastly, inspired by (Yang et al., 2023a; Jiang et al., 2024), we investigate how the volume of training data impacts model performance by conducting a comprehensive experiment in which the training set size ranges from 1% to 100% to examine how well models generalize with few domain samples. We evaluate drug recommendation on MIMIC-IV, since its complexity poses a challenging setting for prediction under varying EHR data sizes. All reported metrics are averaged over five independent runs. The results in Figure 5 indicate that all models show reduced performance in both AUPRC and F1-score when labeled data are scarce, particularly below 10% of the training set. However, UdonCare maintains a considerable edge over other baselines, suggesting that its co-training strategies effectively minimize information loss even when domain features are limited. Notably, MLDG lacks a certain level of resilience against data limitation, likely due to unique domain assignment, which might not work for the situation that existing patients with only few admissions (less than 3) on EHR datasets.
6 Conclusion
In this work, we propose UdonCare, a novel framework for clinical domain generalization that discovers hierarchy-grounded domains with adaptive granularity. By pruning hierarchical concepts, our model addresses two unique challenges: the absence of domain IDs and the need for domain definitions that align with medical semantics. Extensive evaluations on multiple EHR datasets demonstrate that UdonCare outperforms existing baselines across diverse tasks, without incurring additional computational overhead. For example, in mortality prediction, UdonCare achieves a relative improvement of 11% in AUPRC over baselines. In the future, we plan to study multi-modal clinical prediction and more general ontology-rich problems.
Impact Statement
This paper presents work whose goal is to advance the field of domain adaptation in clinical prediction. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here. However, we extend the discussion of some potential directions beyond healthcare event prediction in Appendix L.
Appendix A Related Work: Hierarchy-Aware Predictive Modeling
EHR prediction models can be broadly categorized into three types: RNN/CNN-based (Choi et al., 2017; Tong et al., 2025a), GNN-based (Lu et al., 2021; Jiang et al., 2024), and Transformer-based models (Shang et al., 2019a; Poulain and Beheshti, 2024; Tong et al., 2025b). Concretely, most graph-based studies (Choi et al., 2017; Shang et al., 2019a; Lu et al., 2021) utilize hierarchical medical classifications like ICD-9 (Organization and others, 1988) and ATC (Nahler and Nahler, 2009) to determine medical concept similarity by assuming diseases closer in the hierarchy share more characteristics, as reflected in similar embeddings. However, this method can be biased as it typically fails to capture complex relationships beyond simple parent-child links, such as complications or comorbidity, leading to sub-optimal predictions (Xu et al., 2023, 2024; Hu et al., 2024). Moreover, there is little research integrating these hierarchical structures with DG techniques, which could enhance model robustness across diverse healthcare settings. Properly leveraging hierarchical relationships in DG could improve the domain discovery process, ensuring models account for variance in disease manifestation across different patient demographics and regional practices. Thus, integrating hierarchy-aware modeling with DG approaches holds potential for developing more accurate and personalized predictive models in EHR, catering to the nuanced needs of global healthcare environments.
Appendix B Pseudo Code for UdonCare
Since the training and inference phase has been explained in the main paper, we conduct our pseudo code by two consecutive phases, as shown in Algorithm 1:
Algorithm 1 Overview of UdonCare
1:Input: EHR dataset $\mathcal{S}$ with patient’s data $\{\mathbf{x}^{(i)}\}_{i=1}^{N_{\mathrm{tr}}}$; Feature extractor $f_{\phi}$ with defined backbone (e.g. Transformer).
2:// When iteration $I=3$ and epochs $N=100$
3:for epoch $\in\{1,2,\dots,40\}$ do
4:  // Backbone Pathway
5:  Decode $\hat{y}_{p}\leftarrow d_{\eta}(f_{\phi,k}(x))$ with Equation (4.3);
6:end for
7: Obtain learned $\phi$; Initialize the hierarchy $\mathcal{H}$;
8:for iteration $\in\{1,2,3\}$ do
9:  // Hierarchy-Guided Domain Discovery
10:  Define & Update look-up table $\mathbf{M}$ via Step 1-4;
11:  Assign domain IDs $\mathcal{M}:\mathbf{x}\mapsto\mathbf{m}$ with Equation (2);
12:  // Self-Supervised Domain Encoding
13:  Initialize $g_{\theta}$; Pretrain $g_{\theta}$ by minimizing Equation (6);
14:  for epoch $\in\{1,2,\dots,20\}$ do
15:   // Domain Pathway
16:   Obtain patient embeddings $\mathbf{p}$ with Equation (1);
17:   Get domain features $\mathbf{r}$ by $g_{\theta}(\mathbf{m})$;
18:   Decompose invariant features $\mathbf{h}$ with Equation (7);
19:   Decode $\hat{y}_{p},\hat{y}_{h}$ with Equation (4.3);
20:   Minimize co-training loss $\mathcal{L}_{p},\mathcal{L}_{h}$ with Equation (4.3);
21:  end for
22:end for
23:Output: Trained models $f_{\phi}$, $g_{\theta}$, $d_{\eta}$, $q_{\xi}$; final prediction $\hat{y}_{h}$.
Appendix C Notation Table
For clarity, we have attached a notation table here, summarizing all symbols used in the main paper, as shown in Table 4. The paper uses the generic notation $\mathbf{x}$ to denote patient input features. We use plain letters for scalars (e.g., $d,t,y$), boldface lowercase letters for vectors (e.g., $\mathbf{p},\mathbf{r},\mathbf{h}$), and boldface uppercase letters for matrices (e.g., $\mathbf{M}$). Euler script letters (e.g., $\mathcal{H},\mathcal{C}$) denote sets or hierarchies.
Table 4: Notations used in UdonCare.
Symbols Descriptions
$x^{(i)}$, $V_{t}^{(i)}$ EHR data of the $i$-th patient; the $t$-th visit of the $i$-th patient
$\mathbf{x}=\sum_{t=1}^{T}x_{t}$ Aggregated multi-visit representation used for domain discovery
$\mathcal{C}=\{c_{1},\ldots,c_{|\mathcal{C}|}\}$ Set of medical concepts (leaf-level ICD-9-CM codes)
$y^{(i)}\in\{0,1\}^{d}$ Ground-truth label vector for predictive tasks (binary or multi-label)
$f_{\phi,k}(\cdot)$, $\mathbf{p}_{k}$ Feature encoder for concept type $k$ (diagnosis, treatment, medication)
$\mathbf{p}_{k}=f_{\phi,k}(x_{k})$ The patient embedding for key $k$
$\mathbf{p}=\mathbf{p}_{1}\oplus\dots\oplus\mathbf{p}_{K}$ Concatenated patient-level embedding
$\mathcal{H}$ Concept hierarchy with $H$ levels (ICD-9-CM)
$n_{i}^{(h)}$ Node $i$ at hierarchy level $h$
$\mathrm{Desc}(n)$ Descendant leaf nodes of node $n$
$e_{i}$, $e_{i}^{(h)}\in\mathbb{R}^{h}$ Embedding of concept node or intermediate ontology node
$m:c_{i}\mapsto n_{i}^{(H)}$ Mapping from a medical concept to its leaf-level ontology node
$S(n)$ Score for pruning node $n$ based on purity, coverage, and depth
$\mathrm{cov}(n)$, $\mathrm{pur}(n)$, $\mathrm{dep}(n)$ Coverage, purity, and depth used in node scoring
$\alpha$, $\rho$ Hyperparameters controlling score balance and similarity threshold
$\mathbf{M}\in\{0,1\}^{N_{\mathrm{tr}}\times|C^{\prime}|}$ Domain-assignment matrix after pruning; encodes soft domain labels
$C^{\prime}$ Pruned code vocabulary representing discovered domains
$g_{\theta}(\cdot)$, $\mathbf{r}=g_{\theta}(\mathbf{m})$ Domain encoder producing latent domain representation $\mathbf{r}$
$\bar{\mathbf{p}}=\mathbb{E}[\mathbf{p}|\mathbf{m}]$ Mean patient embedding per domain, used for self-supervised alignment
$L_{r}$ Self-supervised loss (MSE + MMD) for domain encoder training
$\mathbf{h}=\mathbf{p}-\tilde{\mathbf{r}}$ Invariant feature obtained by removing domain-parallel components
$\tilde{\mathbf{r}}=\mathbf{r}\cdot\langle\mathbf{p}/\|\mathbf{r}\|,\,\mathbf{r}/\|\mathbf{r}\|\rangle$ Parallel projection of $\mathbf{p}$ onto domain direction $\mathbf{r}$
$d_{\eta}(\cdot)$, $\hat{y}_{p}$ Backbone prediction head producing $\hat{y}_{p}=d_{\eta}(\mathbf{p})$
$q_{\xi}(\cdot)$, $\hat{y}_{h}$ Invariant prediction head producing $\hat{y}_{h}=q_{\xi}(\mathbf{h})$
$\tilde{y}$ Averaged pseudo-label from $(\hat{y}_{p},\hat{y}_{h})$ for KL consistency
$L_{p}$, $L_{h}$ Mutual learning objectives combining CE loss and KL divergence
$\lambda$ KL divergence weight.
$\langle\cdot,\cdot\rangle$ Vector inner product
$\|\cdot\|$ Vector norm (Euclidean unless specified)
$W_{\text{cls}}(\cdot)$ the learned weight of a logistic regression task
Appendix D Baselines
Beyond Base and Oracle, we select 5 approaches following general DG setting and select 2 highly related baselines tackling clinical DG problems to compare the performance with UdonCare:
  • Domain-adversarial neural networks (DANN) (Ganin et al., 2016) use gradient reversal layer for domain adaptation, and we adopt it for the generalization setting by letting the discriminator only predict training domains.
  • Conditional adversarial net (CondAdv) (Isola et al., 2017) concatenates the label probability and the feature embedding to predict domains in an adversarial way.
  • Meta-learning for domain generalization (MLDG) (Li et al., 2018a) adopts the model-agnostic meta learning (MAML) (Finn et al., 2017) framework for domain generalization.
  • Invariant risk minimization (IRM) (Arjovsky et al., 2019) learns domain invariant features by regularizing squared gradient norm.
  • Proxy-based contrastive learning (PCL) (Yao et al., 2022) build a new supervised contrastive loss from class proxies and negative samples.
  • Many-domain generalization for healthcare (ManyDG) (Yang et al., 2023a) with auto-encoder structures to learn invariant features with unique domain separation for each patient.
  • Self-learning framework for domain generalization (SLDG) (Wu et al., 2023) discovers latent domains by decoupled domain-specific classifiers for clinical prediction.
Note that, for baselines that rely on domain IDs, we use admission time as the domain definition.
Appendix E Data Statistics
Table 5: Statistics of MIMIC-III and MIMIC-IV datasets.
Dataset # patients Max. # visit Avg. # visit Avg. $\mathcal{D}$/visit Avg. $\mathcal{P}$/visit Avg. $\mathcal{M}$/visit
MIMIC-III 6,497 42 2.66 13.06 4.54 33.71
MIMIC-IV 49,558 55 3.66 13.38 4.70 43.89
MIMIC-III and MIMIC-IV are widely used electronic health records in clinical prediction (Johnson et al., 2016, 2023). MIMIC-III covers ICU admissions from 2001 to 2012, while MIMIC-IV spans 2008 to 2019. To avoid overlapping time ranges with MIMIC-III, we only retain patients from the years 2013201320192019 in MIMIC-IV.
Appendix F Implementation Details
Considering the common use of the Encoder-Decoder structure for clinical prediction, we adopt the Transformer (Vaswani et al., 2023) as the backbone feature extractor $f_{\phi}$ in UdonCare and all baselines. Specifically, we follow the implementation adapted from PyHealth (Yang et al., 2023b), consisting of three layers with a hidden size of 64, 4 attention heads, and a dropout rate of 0.2. The position encoding is applied across patient visits to capture temporal order. Diagnosis, treatment, and medication codes are embedded as separate feature keys using an embedding look-up table. The MLP classifiers used for both original and domain-invariant representations contain two hidden layers with sizes [64, 32], ReLU activations, and a dropout rate of 0.2, with task-specific output activations. We use the Adam optimizer for training, and all remaining hyperparameters follow the settings in PyHealth. For domain discovery, the score function uses $\alpha=0.5$, and the KL divergence loss coefficient $\lambda$ is set to 1.0 on MIMIC-III and 1.5 on MIMIC-IV. All models are trained for 100 epochs, and the best model is selected based on the AUPRC score monitored on the source validation set. We set the learning rate to $1\times 10^{-4}$ for $f_{\phi}$ and $5\times 10^{-5}$ for $g_{\theta}$, batch size to 32, iteration to 3, and self-supervised epoch to 30. We tune $\alpha$ and $\lambda$ based on validation performance, which are 5e35e-3 and $0.5$, respectively. All experiments are conducted using Python 3.10 and PyTorch 2.3.1 with CUDA 12.4 on a server equipped with AMD EPYC 9254 24-Core Processors and NVIDIA L40S GPUs.
Appendix G Additional Experimental Results
G.1 Results with GAMENet Backbone
We extend the main experiments by replacing the Transformer backbone with GAMENet (Shang et al., 2019b), a model specifically designed for drug recommendation tasks. Therefore, we evaluate its performance on the drug recommendation task using both MIMIC-III and MIMIC-IV. As shown in Table 6, the gap between Oracle and Base again highlights the domain shift between datasets. Standard domain generalization methods (e.g., DANN, IRM) provide limited improvements, while adversarial and meta-learning approaches (CondAdv, MLDG) show moderate gains. ManyDG performs well on MIMIC-IV, but our proposed UdonCare achieves the best overall results, with consistent improvements in both AUPRC and F1-score. These findings confirm that UdonCare generalizes effectively even with a domain-specific backbone.
Table 6: Performance of Drug Recommendation with GAMENet.
Model MIMIC-III MIMIC-IV
AUPRC F1-score AUPRC F1-score
Oracle 79.57 (0.32) 68.14 (0.27) 75.92 (0.27) 64.57 (0.36)
Base 72.49 (0.17) 58.32 (0.21) 67.88 (0.30) 58.26 (0.27)
DANN 74.24 (0.09) 59.21 (0.17) 72.18 (0.13) 60.35 (0.18)
CondAdv 73.38 (0.11) 62.94 (0.13) 70.86 (0.28) 59.55 (0.27)
MLDG 75.26 (0.14) 63.33 (0.16) 71.49 (0.19) 61.19 (0.38)
IRM 72.13 (0.21) 59.87 (0.25) 71.35 (0.20) 60.87 (0.34)
ManyDG 76.84 (0.10) 64.58 (0.30) 74.18 (0.24) 62.23 (0.29)
UdonCare 77.56 (0.15) 65.79 (0.24) 74.83 (0.21) 61.94 (0.32)
G.2 Results with CGL Backbone
We also extend the main experiments by replacing the Transformer backbone with CGL (Lu et al., 2021), a model specifically designed for diagnosis prediction tasks. Following the original CGL setting, we only consider conditions as input features, which naturally leads to some performance degradation compared to the main experiment results. Therefore, we evaluate its performance on the diagnosis prediction task using both MIMIC-III and MIMIC-IV. As shown in Table 7, the gap between Oracle and Base again highlights the domain shift between datasets. Standard domain generalization methods (e.g., DANN, IRM) provide limited improvements, while adversarial and meta-learning approaches (CondAdv, MLDG) show moderate gains. ManyDG performs competitively, but our proposed UdonCare still achieves the best overall results, with consistent improvements across both w-F1 and R@10. These findings confirm that UdonCare generalizes effectively even with a task-specific backbone.
Table 7: Performance of Diagnosis Prediction with CGL.
Model MIMIC-III MIMIC-IV
w-F1 R@10 w-F1 R@10
Oracle 25.71 (0.18) 38.01 (0.21) 27.09 (0.22) 39.42 (0.19)
Base 19.05 (0.14) 30.01 (0.24) 19.22 (0.25) 30.64 (0.18)
DANN 20.35 (0.20) 33.56 (0.27) 23.02 (0.15) 34.07 (0.23)
CondAdv 21.55 (0.13) 35.11 (0.22) 25.02 (0.28) 36.51 (0.16)
MLDG 20.12 (0.26) 32.81 (0.19) 23.15 (0.19) 33.27 (0.21)
IRM 21.02 (0.12) 32.11 (0.25) 22.41 (0.27) 33.01 (0.22)
ManyDG 22.01 (0.23) 35.11 (0.20) 24.12 (0.17) 36.02 (0.29)
UdonCare 23.89 (0.20) 37.12 (0.24) 26.02 (0.15) 38.21 (0.18)
G.3 Results of Cross-Institutional Distributional Shift
Following the setting used in SLDG (Wu et al., 2023), we extend our evaluation to the eICU dataset, where hospitals can be categorized into four groups based on their locations (Midwest, Northeast, West, and South). To quickly deploy our model on this dataset, we use the processing procedure of PyHealth. We evaluate the performance of our model under spatial domain shifts by treating the Midwest group as the target domain and the remaining groups as the source domain. We use diagnosis and treatment codes to construct patient records, as medications in eICU are not encoded with standard IDs. Table 8 reports the results for readmission and diagnosis prediction. We observe that UdonCare continues to demonstrate the superiority over all baselines across these two tasks.
Table 8: Performance under Cross-Institutional Distributional Shift on eICU.
Model Readmission Prediction Diagnosis Prediction
AUPRC AUROC w-F1 R@10
Oracle 22.74 (0.13) 69.29 (0.13) 63.18 (0.08) 79.05 (0.09)
Base 11.97 (0.08) 51.31 (0.06) 53.89 (0.05) 71.40 (0.05)
DANN 12.75 (0.16) 54.28 (0.14) 58.41 (0.18) 75.66 (0.24)
CondAdv 13.16 (0.18) 53.97 (0.12) 57.04 (0.23) 73.92 (0.17)
MLDG 14.82 (0.11) 54.25 (0.17) 58.13 (0.28) 75.31 (0.21)
IRM 15.73 (0.19) 58.31 (0.21) 59.22 (0.16) 76.46 (0.15)
PCL 15.14 (0.23) 58.06 (0.24) - -
ManyDG 17.53 (0.17) 60.20 (0.10) 60.61 (0.38) 76.94 (0.21)
SLDG 16.71 (0.21) 59.54 (0.18) - -
UdonCare 18.37 (0.18) 62.12 (0.11) 61.83 (0.14) 77.02 (0.11)
Appendix H Domain Discovery with More Feature Keys
To examine whether latent domain discovery benefits from richer feature information, we further incorporate procedure and medication codes as additional keys in UdonCare. As shown in Table 9, the overall gains across prediction tasks are marginal and vary inconsistently across datasets. While certain metrics observe slight improvements, others remain unchanged or even decline. This suggests that introducing conditions, treatments and drugs together into the domain partitioning process does not lead to stable enhancements. These findings are also consistent with our understanding, since we are more likely to get fine-grained domains: patients with the same disease may be split into different domains solely due to distinct medication regimens, which can weaken the benefits of ontology-guided grouping. Therefore, our conclusion is simply that using several ontologies at the same time is less efficient. Instead, we encourage practitioners to select the most appropriate ontology according to the task.
Note that, our reliance on the ICD-9-CM is not due to a limitation of our model, but because diseases are the most direct and required source of patient information. Almost all EHR prediction tasks rely on diagnostic codes as a primary input, and diseases exert strong influence on domain formation. Hence, the disease ontology serves as the most appropriate basis in the general clinical setting. It is also important to note that UdonCare is not restricted to disease ontologies. For instance, using a medication ontology (ATC classification) might be equally or even more appropriate in drug recommendation. Our model directly supports such substitution. In this paper, we want to align the used ontology across four tasks in our experiments.
Table 9: Performance comparison of four prediction tasks on MIMIC-III/MIMIC-IV. We report the average performance (%) and the standard deviation (in bracket) over 5 runs.
Model Task 1: Mortality Prediction Task 2: Readmission Prediction
MIMIC-III MIMIC-IV MIMIC-III MIMIC-IV
AUPRC AUROC AUPRC AUROC AUPRC AUROC AUPRC AUROC
Oracle 16.95 (0.48) 70.61 (0.58) 8.46 (0.53) 68.92 (0.47) 74.42 (0.43) 69.74 (0.48) 69.37 (0.12) 68.47 (0.11)
Base 11.05 (0.65) 55.01 (0.98) 3.97 (0.47) 59.13 (0.76) 63.45 (0.84) 62.15 (0.73) 58.46 (0.24) 60.56 (0.31)
UdonCare 16.78 (0.35) 70.02 (0.40) 7.21 (0.29) 67.18 (0.46) 72.31 (0.33) 68.32 (0.36) 62.65 (0.11) 59.14 (0.23)
Model Task 3: Drug Recommendation Task 4: Diagnosis Prediction
MIMIC-III MIMIC-IV MIMIC-III MIMIC-IV
AUPRC F1-score AUPRC F1-score w-F1 R@10 w-F1 R@10
Oracle 80.43 (0.14) 67.41 (0.29) 74.31 (0.25) 61.28 (0.22) 26.97 (0.13) 39.46 (0.27) 28.12 (0.11) 40.53 (0.16)
Base 68.83 (0.19) 57.31 (0.35) 66.94 (0.18) 53.13 (0.18) 21.32 (0.19) 31.08 (0.23) 20.07 (0.12) 31.52 (0.18)
UdonCare 79.34 (0.20) 66.98 (0.30) 73.68 (0.35) 60.22 (0.15) 25.28 (0.12) 38.97 (0.21) 27.86 (0.12) 40.41 (0.18)
Appendix I Convergence Analysis via Iterative Learning
While the pruning-based algorithm provides efficiency, its iterative nature makes it non-trivial to characterize convergence using a simple continuous optimization view. To further substantiate the convergence property of our approach, we extend the cosine similarity experiment described in Section 5.4. Specifically, instead of reporting a single snapshot after three iterations, we monitor the cosine similarity values across iterations. As shown in Table 10, the results exhibit a clear trend: at the early stage of training (epoch 40), the similarities between $W_{d}(\mathbf{p}\rightarrow\text{labels})$ and $W_{d}(\mathbf{h}\rightarrow\text{labels})$ are relatively low, while the cross-domain similarities ($W_{d}(\mathbf{p}\rightarrow\text{domains})$ vs. $W_{d}(\widetilde{\mathbf{r}}\rightarrow\text{domains})$) are comparatively high. This indicates that the model has not yet disentangled domain- and label-related features. As the training proceeds (epoch 60 and 80), the similarities gradually align with the final values at epoch 100, where the decomposition becomes stable and consistent with the results reported in Table 2. These observations provide additional evidence that the iterative learning strategy enables the model to converge in terms of separating label-invariant and domain-specific information. More importantly, by tracking cosine similarity dynamics, we validate that pruning and iterative decomposition jointly lead to a stable representation space, rather than an artifact of a single training snapshot.
Table 10: Cosine similarity of linear weights on $\mathbf{p}$, $\widetilde{\mathbf{r}}$, and $\mathbf{h}$ across epochs.
Epoch Cosine Similarity Mortality Prediction Readmission Prediction Drug Recommendation Diagnosis Prediction
40 $W_{d}(\mathbf{p}\rightarrow\text{labels})$ vs. $W_{d}(\mathbf{h}\rightarrow\text{labels})$ 0.5283 0.2935 0.3829 0.6259
$W_{d}(\mathbf{p}\rightarrow\text{domains})$ vs. $W_{d}(\widetilde{\mathbf{r}}\rightarrow\text{domains})$ 0.4615 0.4908 0.5275 0.4705
$W_{d}(\mathbf{p}\rightarrow\text{labels})$ vs. $W_{d}(\mathbf{p}\rightarrow\text{domains})$ 0.2495 0.2695 0.2993 0.3826
60 $W_{d}(\mathbf{p}\rightarrow\text{labels})$ vs. $W_{d}(\mathbf{h}\rightarrow\text{labels})$ 0.7245 0.4119 0.6080 0.8318
$W_{d}(\mathbf{p}\rightarrow\text{domains})$ vs. $W_{d}(\widetilde{\mathbf{r}}\rightarrow\text{domains})$ 0.3960 0.2195 0.3450 0.2712
$W_{d}(\mathbf{p}\rightarrow\text{labels})$ vs. $W_{d}(\mathbf{p}\rightarrow\text{domains})$ 0.1842 0.1007 0.1812 0.1437
80 $W_{d}(\mathbf{p}\rightarrow\text{labels})$ vs. $W_{d}(\mathbf{h}\rightarrow\text{labels})$ 0.7761 0.4686 0.6392 0.8523
$W_{d}(\mathbf{p}\rightarrow\text{domains})$ vs. $W_{d}(\widetilde{\mathbf{r}}\rightarrow\text{domains})$ 0.3508 0.2012 0.2807 0.2208
$W_{d}(\mathbf{p}\rightarrow\text{labels})$ vs. $W_{d}(\mathbf{p}\rightarrow\text{domains})$ 0.1312 0.0853 0.1296 0.1017
100 $W_{d}(\mathbf{p}\rightarrow\text{labels})$ vs. $W_{d}(\mathbf{h}\rightarrow\text{labels})$ 0.7831 0.4813 0.6546 0.8654
$W_{d}(\mathbf{p}\rightarrow\text{domains})$ vs. $W_{d}(\widetilde{\mathbf{r}}\rightarrow\text{domains})$ 0.3427 0.1957 0.2753 0.2133
$W_{d}(\mathbf{p}\rightarrow\text{labels})$ vs. $W_{d}(\mathbf{p}\rightarrow\text{domains})$ 0.1239 0.0794 0.1251 0.0972
* $W_{d}(\cdot)$ represents the learned linear weights. Cosine similarity scores are averaged across all classes and evaluated over 3 runs.
Appendix J Upward Information Flow
To enable hierarchy-aware domain discovery, the model must obtain embeddings for all nodes in the ontology $\mathcal{H}$, not only the leaf concepts $\mathcal{C}$. This appendix details the complete procedure for (i) initializing node features, (ii) propagating them upward through the hierarchy, and (iii) refining ancestor embeddings using hierarchical similarity.
Initializing Leaf-Node Embeddings First, for each diagnostic code $d_{i}$, we can initialize its embedding $\mathbf{e}_{d_{i}}\in\mathbb{R}^{h}$ in terms of one of two scenarios:
  1. 1.
    For present code $d_{i}$ in the dataset, $\mathbf{e}_{d_{i}}$ is initialized by embedding table in disease-specific extractor $f_{\phi,d}(\cdot)$.
  2. 2.
    For absent code $d_{i}$ in the dataset, $\mathbf{e}_{d_{i}}$ is initialized by its entity name through the pretrained ClinicalBERT (Huang et al., 2019).
Bottom-Up Propagation in the Ontology Since internal nodes have no direct representations, we propagate information from leaf nodes upward. For each parent node $n_{i}^{(h-1)}$ at level h1h-1, its embedding is initialized by averaging the embeddings of its descendants at level $h$:
$\mathbf{e}^{(h-1)}_{i}\;=\;\frac{1}{|\mathrm{Desc}(n^{(h-1)}_{i})|}\sum_{n\in\mathrm{Desc}(n_{i})}\mathbf{e}^{(h)}_{n}$,
which extends the leaf-level table $\text{E}(e_{1},\dots,e_{|\mathcal{C}|})$ to $\text{E}(e_{1},\dots,e_{|\mathcal{H}|})$. However, these initial embeddings do not capture the hierarchical distances between node pairs. For example, two codes might have significantly different embeddings despite sharing the same parent node.
LCA-Based Refinement via Hierarchical Similarity Averaging descendants does not reflect hierarchical distances: two nodes may have similar features even if their parents are far apart in $\mathcal{H}$. To encode such structure, we incorporate a hierarchical-clustering refinement (Johnson, 1967).
(1) Compute pairwise cosine similarity. For any two nodes $(d_{i},d_{j})$ (leaf nodes only), we compute
$\text{cos}(\mathbf{e}_{d_{i}},\mathbf{e}_{d_{j}})=\frac{\mathbf{e}_{d_{i}}^{\top}\mathbf{e}_{d_{j}}}{\|\mathbf{e}_{d_{i}}\|\,\|\mathbf{e}_{d_{j}}\|}.$
Among all possible pairs, we select the one with the highest cosine similarity.
(2) Update the lowest common ancestor (LCA). Let $n_{\mathrm{LCA}}(d_{i},d_{j})$ denote their lowest common ancestor. We update its embedding by averaging itself with the embeddings of the two nodes:
$\mathbf{e}_{\mathrm{LCA}(d_{i},d_{j})}\leftarrow\mathrm{Average}\!\left(\mathbf{e}_{\mathrm{LCA}(d_{i},d_{j})},\mathbf{e}_{d_{i}},\mathbf{e}_{d_{j}}\right).$
(3) Iterative refinement. We repeat this process along the ordered similarity list and stop once the maximum remaining similarity falls below a predefined threshold $\rho=0.3$. Nodes never selected as LCAs retain their initial average-based embeddings from Eq. (3).
Appendix K Scoring Function
The candidate score $S(n)$ involves three indicators:
  1. 1.
    Coverage, $\mathrm{cov}(n)$: measures the ratio of covered leaf nodes $n\in\mathcal{L}$. A higher value indicates greater coverage, favoring the selection of higher-level nodes:
    $\mathrm{cov}(n)=\frac{|\mathcal{M}|}{|\mathcal{L}|},$
    where $\mathcal{M}=\mathrm{Desc}(n)$ denotes the set of descendant leaf nodes of $n$, and $|\mathcal{L}|$ is the total number of leaf nodes in the hierarchy. A higher value indicates that $n$ covers a larger portion of the tree, thus favoring higher-level nodes.
  2. 2.
    Purity, $\mathrm{pur}(n)$: gauges how cohesive the partitioned tree is in the embedding space. This metric evaluates how coherent the subtree rooted at $n$ is in the embedding space. It is computed as the expected similarity between the embedding of node $n$ and the embeddings of its descendant leaf nodes:
    $\mathrm{pur}(n)=\mathbb{E}_{m\in\mathcal{M}}\left[\mathrm{sim}(\mathbf{e}_{n},\mathbf{e}_{m})\right],$
    where $\mathrm{sim}(\cdot)$ denotes the similarity function (e.g., cosine similarity). Higher purity implies that leaf nodes under $n$ are more semantically aligned with the root node.
  3. 3.
    Depth, $\mathrm{dep}(n)\in\{1,\dots,H\}$: indicates the hierarchical level of node $n$. It is defined as:
    $\mathrm{dep}(n)=\frac{h}{H},$
    where $h$ is the depth of node $n$ and $H$ is the maximum depth of the hierarchy. A larger value corresponds to finer-grained concepts, thus favoring deeper nodes.
Appendix L Broader Applicability and Extensions of UdonCare
While UdonCare is evaluated primarily on structured, code-based EHR data, the core design of the framework is modality-agnostic and can be extended beyond the four clinical event prediction tasks explored in this work. We briefly highlight two promising directions:
L.1 Extension to Multi-modal Clinical Data
The ontology-guided domain discovery mechanism is compatible with richer, multi-modal patient records. Modern hospital systems routinely collect diverse modalities, including:
  1. 1.
    Clinical notes (text), encoded with contextualized language models;
  2. 2.
    Radiology and pathology images, structured under well-defined ontologies such as RadLex;
  3. 3.
    Physiological waveforms (ECG, EEG), which contain temporally structured signals;
  4. 4.
    Laboratory trajectories, which follow standardized concept hierarchies (LOINC).
In principle, our model can incorporate these modalities by (1) extending the feature extractor to relevant encoders (e.g., CNNs, transformers for images, waveform encoders), and (2) pruning cross-modal ontologies or multi-level taxonomies whenever available. The mutual-learning mechanism in UdonCare, designed to decouple domain and label information, benefits from such multi-modal richness by preventing modality-specific biases from dominating the latent domain assignments. However, large-scale multi-modal extensions require access to paired datasets, which is unavailable in MIMIC or other standard public EHR databases. Thus, multi-modal generalization remains an important but data-dependent direction for future research.
L.2 Generalization Beyond Healthcare to Ontology-Rich Domains
The ontology-guided design is also not restricted to clinical settings. Many real-world domains exhibit hierarchical or taxonomy-driven feature structures that mirror the ICD hierarchy, such as
  1. 1.
    E-commerce: product category graphs, brand hierarchies, user segments;
  2. 2.
    Recommendation systems: taxonomies of items or content genres;
  3. 3.
    Knowledge-driven applications such as academic topic hierarchies, legal codes, or biological ontologies.
In these domains, domain shifts commonly arise from evolving user groups, temporal drifts in product popularity or region-specific differences, which are analogous to the clinical temporal and spatial shifts studied in this work. The pruning-based latent domain discovery can meaningfully identify stable high-level clusters across such hierarchies, while the invariant-projection mechanism can separate domain-specific variations from task-relevant signals. The adaptability of its ontology-guided components suggests broad potential applicability to many hierarchical, knowledge-grounded machine learning scenarios.
Appendix M More Explanations on the Learned Linear Weights
In this section, we discuss the relatively low cosine similarity in Table 2:
First, a similar trend has been observed in ManyDG (Yang et al., 2023a). Linear probing for domain prediction is inherently more challenging than probing for label prediction. The domain classification task usually contains a substantially larger number of categories (e.g. over 500), whereas the label space is considerably smaller. Thus, the learned linear coefficients for domain prediction tend to be less aligned, resulting in lower cosine similarity in the second row.
Second, if the model successfully decomposes patient embeddings into label-related and domain-related components, the observed similarity pattern becomes intuitive. The first row corresponds to label prediction, and thus naturally shows higher similarity because label information constitutes a larger and stronger portion of the embedding space. In contrast, after label (domain-invariant) information is accounted for, the remaining domain signal becomes weaker, making it harder for a linear classifier to capture consistently. This yields lower similarity in the second row.
We emphasize that this explanation does not imply a precise quantitative ratio between label and domain information. Rather, the key observation is that when the invariant (label-relevant) component dominates the embedding space, the remaining domain-specific signal becomes harder to recover linearly, and this phenomenon is reflected consistently in our empirical results.
Appendix N The Use of Large Language Models (LLMs)
In preparing this paper, we used large language models (LLMs) solely as a general-purpose tool to improve writing fluency and polish the presentation of the text. All ideas, experimental designs, analyses, and conclusions are our own, and the responsibility for the content rests entirely with the authors.